# Time-Spectral Resolvent AnalysisM. Howell and S. He

### v1

@@ 1 Introduction
The classical resolvent analysis was formulated as a tool to identify the most amplified output structures of fluid flows when subject to a harmonic forcing. Amplification analysis of fluid flows with harmonic forcing was first presented by Trefethen1993 as a framework to show how linearly stable flows can still have large amplifications when subject to harmonic perturbations. The idea that specific output modes can be amplified by different perturbations was further explored by Bamieh2001 who discovered that even in subcritical conditions, the Navier-Stokes operator can amplify inputs and cause large response gains. JOVANOVI2005 later explored the exact input modes that caused the most amplified responses in the velocity field for channel flows. The idea of resolvent analysis was then formalized in the context of wall-bounded turbulence by McKEON2010, who showed that dominant flow features could be understood as the most amplified resolvent modes. The resolvent operator maps between an external harmonic forcing and a deviation from a steady base flow and the singular value decomposition (SVD) of the resolvent operator identifies the most amplified input-output structures. Subsequent studies extended the resolvent framework to turbulent channels, boundary layers, and the more advanced transonic buffet [hwang2010linear, Sipp2012, Kojima2019]. Recent reviews [taira2017modal, McKeon2017] have established resolvent analysis as a central tool for connecting linear amplification mechanisms with nonlinear flow dynamics.
@@
Extending resolvent analysis to flows with time-periodic base states has posed significant challenges such as quasi-periodic responses to forcing and the computational expense of linearizing around the entire base flow. Several approaches have been developed to address these challenges, including mean flow analysis, Koopman-based methods, and harmonic resolvent formulations that explicitly account for frequency interactions.
The simplest approach is mean flow resolvent analysis, which uses the time-averaged base flow to predict the resolvent amplification. This idea was analyzed in Turton2015 and Leclercq2023, where it was found that mean flow analysis is only valid if the resolvent operator of the mean flow exhibits a single dominant singular value (i.e $\sigma_{1}\gg\sigma_{2}$). Also, the flow must exhibit quasi-monochromatic behavior as the interaction of higher harmonics in the flow can disrupt the stability of the mean flow. This idea was later extended to cases where the periodic base flow has a well-defined statistically steady state in [Leclercq2023]. In this case, the flow is first linearized over the statistically steady base flow, then the phase averaged resolvent operator is computed. This extends resolvent analysis to a more general operating point; however, it does not take into account the cross-frequency interactions between the base flow and harmonic forcing.
@@
Recent theoretical work by Susuki2021 has extended Laplace domain analysis to nonlinear limit cycles using the Koopman resolvent. This operator is the Laplace realization of the Koopman generator [Koopman1931] and decomposes the resolvent spectrum into stationary modes which correspond to base-flow harmonics and nonstationary modes which are governed by Floquet exponents. However, the rigorous application of such resolvent operators to quasi-periodic or rotating systems is non-trivial; Eiter2022 demonstrated that for irrational frequency ratios, the associated operator may lack a uniform bound due to the accumulation of singularities. These theoretical hurdles motivate the need for a formulation that can robustly capture the dynamics of time-varying systems by explicitly resolving the coupling between the base flow and perturbation harmonics.
The harmonic resolvent formalism explicitly accounts for the full periodic base flow and frequency interactions. Padovan2020 developed this framework to analyze amplification mechanisms and cross-frequency interactions in nonlinear flows, showing how forcing at one frequency can induce responses at other harmonics through quadratic nonlinearity. In later works [Padovan2022, Wu2022, Padovan2024] the harmonic resolvent framework was extended to include sub-harmonics, coupling a forcing frequency with a non-commensurate base frequency. This formulation constructs a resolvent operator with a block-Toeplitz structure that maps the Fourier coefficients of the forcing to the Fourier coefficients of the quasi-periodic response, thereby extending resolvent analysis beyond the single-frequency, steady-base setting. This analysis was later used by Islam2024 to study frequency interactions in compressible cavity flow across an airfoil, where the authors identify how the nonlinear flow permits energy transfer across frequencies in a flow with multiple resonant frequencies. However, the frequency-domain formulation requires computing Fourier coefficients of the periodic Jacobian and assembling block-Toeplitz operators, which complicates implementation and integration with time-domain solvers.
@@
Beyond the formulation challenges, resolvent analysis for high-dimensional systems also faces computational challenges associated with handling large operators, such as memory allocation. This has been addressed using matrix-free methods [Bagheri2009, MONOKROUSOS2010, Loiseau2018, Martini2021], however these approaches require an additional solver for the adjoint equations. In order to address this challenge in the context of the harmonic resolvent, Farghadan2024 proposed an efficient adjoint-free time-stepping algorithm that avoids explicit formation of the large block-Toeplitz operator, making the method tractable for high-dimensional applications. Other approaches include data-driven resolvent analysis [Herrmann2021] and spectral methods for computing the base flow [Mundis2013].
@@
Rather than directly computing the base flow of a nonlinear system in the time domain using a time-marching solver, spectral methods [Trefethen2000, boyd2013chebyshev] have been shown to be more efficient in solving for the base flow. While spectral methods have traditionally only been used for fully periodic flows, extensions have been made to quasi-periodic cases [Wirth1997] and later shown to scale up to aerospace problems [Mavriplis2011, Mundis2013]. Junge2021 extended spectral methods to cases with multiple non-commensurate frequencies in high dimensional problems by evaluating the frequencies on a torus. In this work, we leverage the accuracy and efficiency of the time-spectral method to extend harmonic resolvent analysis to discrete time. This allows for the exploitation of the computational efficiency of the time-spectral method as well as spectral convergence to the true solution.
@@
This paper is organized as follows: in section 2 we present the classical resolvent analysis for stationary base flows. We then review the frequency domain harmonic resolvent operator from Padovan2020 in section 3. The formulation for the full time-spectral resolvent (TSR) operator using the time-spectral method is presented in section 4 and the special handling and projection required for autonomous systems is detailed in 5. The spectral convergence properties of the TSR and harmonic resolvent operators are compared in section 6. Finally, we validate our TSR operator against direct time accurate integration in section 7 for three examples: a linear system with zero base flow using the Mathieu oscillator, an autonomous non-linear system with the van der Pol oscillator, and a high dimensional autonomous system using the complex Ginzburg-Landau partial differential equation (PDE).

### v2

@@ 1 Introduction
The classical resolvent analysis was formulated as a tool to identify the most amplified output structures of fluid flows when subject to a harmonic forcing. Trefethen1993 first analyzed the linearized Navier-Stokes operator to show that large transient amplifications can still occur in linearly stable flows. It was shown that these amplifications were the result of a linear mechanism caused by the non-normal nature of the Navier-Stokes operator. Bamieh2001 then demonstrated that different input structures can amplify specific output modes and cause large energy amplifications for the Navier-Stokes operator, even in subcritical conditions. The exact input modes which caused the most amplified responses in the velocity field for channel flows were then explored by JOVANOVI2005, showing that the magnitude of amplification depends on the direction of forcing. The resolvent analysis was then formalized in the context of turbulent pipe flow by McKeon2010, who showed that the most amplified input-output structures of a flow could be identified by the resolvent operator. The resolvent operator maps between an external harmonic forcing and a deviation from a steady base flow and the singular value decomposition (SVD) of the resolvent operator identifies the most amplified input-output structures. Subsequent studies applied the resolvent framework to turbulent channels, boundary layers, and the more advanced transonic buffet [hwang2010linear, Sipp2012, Kojima2019]. Recent reviews [taira2017modal] have established resolvent analysis as a central tool for connecting linear amplification mechanisms with nonlinear flow dynamics.
@@
The most basic form of resolvent analysis only considers the response of a linearized system in the vicinity of a steady-state solution. However, it is often desired to analyze systems in the vicinity of an unsteady solution or a time-periodic base flow. Extending resolvent analysis to flows with time-periodic base states has posed significant challenges such as quasi-periodic responses and the computational expense of linearizing around the entire base flow. Also, as described by Leclercq2023, another challenge is that the response of the system is parameterized by the phase of the base flow when the forcing is first applied, making the response of the system to forcing non-unique. Several approaches have been developed to address these challenges including mean flow and mean resolvent analysis, Koopman-based methods, and harmonic resolvent formulations that account for the time-varying nature of the response.
One method to address theses challenges is with the mean resolvent analysis proposed by Leclercq2023, bongarzone2025. As opposed to mean-flow resolvent analysis which first linearizes the dynamics about the time-averaged steady mean flow (as in [Sipp2012]), the mean resolvent analysis first linearizes the dynamics about an ensemble of trajectories on the systems attractor. The mean resolvent operator then predicts the mean response which is statistically optimal in the sense that it is the best linear time invariant (LTI) approximation to the response, thus eliminating the phase dependence of the problem. While the mean resolvent approach accounts for interactions between the perturbations and the unsteady components of the base flow, it filters out the phase-dependent frequency modulation. Consequently, the mean resolvent operator is unable to predict the full quasi-periodic response often desired when analyzing systems with time-periodic base flows.
@@
Recent theoretical work by Susuki2021 has extended Laplace domain analysis to nonlinear limit cycles using the Koopman resolvent. This operator is the Laplace realization of the Koopman generator [Koopman1931], providing a framework to analyze nonlinear autonomous systems without the need for linearization. In the case of time-periodic base flows, the Koopman resolvent decomposes the spectrum into stationary modes which correspond to base-flow harmonics and nonstationary modes which are governed by Floquet exponents. However, the rigorous application of such resolvent operators to quasi-periodic or rotating systems is non-trivial; Eiter2022 demonstrated that for irrational frequency ratios, the associated operator may lack a uniform bound due to the accumulation of singularities. These theoretical hurdles motivate the need for a formulation that can robustly capture the dynamics of time-varying systems by explicitly resolving the coupling between the base flow and perturbation harmonics.
Another approach to predict the exact phase dependent input-output structures for time varying base flows is the harmonic resolvent presented by Padovan2020, Padovan2022. The harmonic resolvent formulation involves linearizing around the exact time-periodic base flow of the system then forming the truncated harmonic resolvent operator by taking the Fourier coefficients of the time-varying Jacobian. This operator has a block-Toeplitz structure and maps between an exponentially modulated periodic (EMP) forcing and response. When comparing the harmonic resolvent framework with that of the mean resolvent analysis [Leclercq2023], it can be noticed that the mean resolvent operator is found along the diagonal block of the harmonic resolvent operator which does not contain any phase-dependent terms. Because the harmonic resolvent maps between EMP signals, it is able to predict the full, phase dependent response, thereby rigorously extending the resolvent framework to time-periodic base flows. However, the frequency-domain formulation requires computing Fourier coefficients of the periodic Jacobian and assembling block-Toeplitz operators, which complicates implementation and integration with time-domain solvers.
@@
Beyond the formulation challenges, resolvent analysis for high-dimensional systems also faces computational challenges associated with handling large operators, such as memory allocation. This has been addressed using matrix-free methods [Bagheri2009, MONOKROUSOS2010, Loiseau2018, Martini2021], however these approaches only consider the mean-flow resolvent. In order to address this challenge in the context of the harmonic resolvent, Farghadan2024 proposed an efficient matrix-free time-stepping algorithm that avoids explicit formation of the large block-Toeplitz operator, making the method tractable for high-dimensional applications. Other approaches include data-driven resolvent analysis [Herrmann2021] and spectral methods for computing the base flow [Mundis2013].
@@
Rather than directly computing the base flow of a nonlinear system in the time domain using a time-marching solver, time-spectral methods [Trefethen2000, boyd2013chebyshev] have been shown to be more efficient in solving for the base flow. While spectral methods have traditionally only been used for fully periodic flows, extensions have been made to quasi-periodic cases [Wirth1997] and later shown to scale up to aerospace problems [Mavriplis2011, Mundis2013]. Junge2021 extended spectral methods to cases with multiple non-commensurate frequencies in high dimensional problems by mapping the quasi-periodic time dynamics onto a multi-dimensional torus. In this work, we leverage the accuracy and efficiency of the time-spectral method to extend harmonic resolvent analysis to discrete time. This allows for the exploitation of the computational efficiency of the time-spectral method as well as spectral convergence to the true solution.
@@
This paper is organized as follows: in section 2 we present the classical resolvent analysis for stationary base flows. We then review the frequency domain harmonic resolvent operator from Padovan2022 in section 3. The formulation for the full time-spectral resolvent (TSR) operator using the time-spectral method is presented in section 4 and the special handling and projection required for autonomous systems is detailed in 5. The spectral convergence properties of the TSR and harmonic resolvent operators are compared in section 6. Finally, we validate our TSR operator against direct time accurate integration in section 7 for three examples: a linear system with zero base flow using the Mathieu oscillator, an autonomous non-linear system with the van der Pol oscillator, and a high dimensional autonomous system using the complex Ginzburg-Landau partial differential equation (PDE).